OpenStudy (anonymous):

3y=-6x+27 y=-2x+12 is this consistent? inconsistent? Dependent?

OpenStudy (anonymous):

can someone please help me? im so lost

OpenStudy (mertsj):

Did you try to solve it?

OpenStudy (mertsj):

Replace y in the first equation with -2x+12 from the second equation and see what you get.

OpenStudy (anonymous):

i got -6x+36=-6x+27

OpenStudy (anonymous):

so it would be inconsistent?

OpenStudy (mertsj):

If you add 6x to both sides, you get 36=27 which is, of course, ridiculous. And it tells you that you are trying to find something that doesn't exist (the point of intersection of 2 parallel lines). So yes, there is no solution and the system is inconsistent.

OpenStudy (anonymous):

okay so how about 6x-2y=18 3x+y=-4

OpenStudy (mertsj):

I would multiply the second equation by 2 and then add the two equations. That would get rid of the y

OpenStudy (anonymous):

so 6x+2y=-8 which would be inconsistent again?

OpenStudy (mertsj):

6x+2y=-8 6x-2y= 18 12x=10 x=5/6

OpenStudy (mertsj):

Did you add the two equations?

OpenStudy (anonymous):

12x+4y=10

OpenStudy (mertsj):